Hello can anyone answer this question ??

In the compound CH2CO2 the hybridisation of the two carbon atoms present are sp2 and sp.
My question is how do u find that it is sp2 and sp??
By seeeing that it is linked with 2,1, or no hydrogen atoms or is there someother way to find the hybridisation??

Help me to solve such questions and plz plz plz ...if u got examples pls do give them to explain it to me a lot better...

Thank you !!!!!!!

Grade:12

2 Answers

vikas askiitian expert
13 years ago

n =[ number of electron in valence shell + number of monovalent atoms attached(like H,Cl etc) - charge on molecule]/2

for n = 2 - hybridisation is SP ,

n =3 hybridisation is SP2 ,

for n = 4 - SP3  & so on  .........

 

see in CH2CO2 we can use above method ,

left carbon bearing two monovalent atoms (2 Hydrogen) so for this carbon n = (4+2)/2 = 3, 

  hybridisation is SP3 ...

right carbon has no monovalent  atoms so n = 4/2 = 2 , its hybridisation is SP ...

 

for hydrocarbons , we have to use another method ...

if carbon is connected to 1 triple bonds  then its hybridisation is Sp ...

if ,,        ,,        ,,           2double bonds then  ,,   ,,       again SP ..

if ,,        ,,        ,,          1 double bond then    ,,     ,,  is SP2

if ,,      ,,           ,,          no double or triple bond ,,  ,,  is SP3

organic compounds do not show hybridisation other than these three ...

 

example - CH2 = C=CH-CH-CH3

from left , 1st carbon - SP2 (only one double bond)

              2nd carbon - SP (two double bonds)

             3rd - SP2

             4rth &  5th -SP3  (no double or triple bond )

ramendra bahadur singh
13 years ago

N=(NO.OF VALENCE ELECTRON + NO. OF SURROUNDING ATOM EXCEPT<O2>+_ CHARGE)  HENCE THE HYBRIDISATION GIVEN IS CORRECT          FOR N=3 SP2     ,       FOR N=2 SP